OpenStudy (anonymous):

What is 9radical 2 minus radical 50

OpenStudy (jdoe0001):

\(\bf 9\sqrt{2}-\sqrt{50}?\)

OpenStudy (anonymous):

Yes

OpenStudy (nfcfox):

Man if u want some long retricedecimal than just leave it as that

OpenStudy (jdoe0001):

\(\bf 9\sqrt{2}-\sqrt{50}\quad \begin{cases} 50\to 2\cdot 5\cdot 5\to 2\cdot 5^2 \end{cases}\quad thus\quad 9\sqrt{2}-\sqrt{2\cdot 5^2} \\ \quad \\ 9\sqrt{2}-5\sqrt{2}\implies ?\)

OpenStudy (anonymous):

Thanks for the answerr

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

7radical 3plus radical 75